OpenStudy (fanduekisses):

f(x) = Sqrt x + seven.; g(x) = 8x - 11 Find f(g(x)) , I keep getting it wrong ;(

OpenStudy (fanduekisses):

I thought it was \[\sqrt{8x-11+7}\]

OpenStudy (anonymous):

yes it is right :) but simplify it more "-11+7"

OpenStudy (fanduekisses):

\[\sqrt{8x-4}\]

OpenStudy (fanduekisses):

but it says the right answer is \[\sqrt[2]{2x-1}\], HOW?

OpenStudy (anonymous):

\(\sqrt{4(2x-1})=2\sqrt{2x-1}\)

OpenStudy (fanduekisses):

oooooooooh so factor then simply the 4 , 2 is its qrt so it goes outside the sqrt

OpenStudy (fanduekisses):

Thanks yall :)